At the instant an electron enters a uniform magnetic field of 0.500 T directed along the positive x-axis, its velocity has the following components: v x =2.00*10^ 6 m / s ,v y =0 and v z =3.00*10^ 6 m / s . Find the radius and pitch of the helical path followed by the electron.


1
Expert's answer
2022-01-10T09:16:07-0500

Given:

"B=0.500\:\rm T"

"v_x=2.00*10^6\:{\rm m/s},\; v_y=0,\;v_z=3.00*10^6\:{\rm m/s}"


The radius of helical path

"R=\frac{mv_z}{eB}=\frac{9.1*10^{-31}*3.00*10^6}{1.6*10^{-19}*0.500}=3.4*10^{-5}\:\rm m"

The period of rotation

"T=\frac{2\pi m}{eB}=\frac{6.28*9.1*10^{-31}}{1.6*10^{-19}*0.500}=7.14*10^{-11}\:\rm s"

The pitch of the helical path

"h=v_xT=2.00*10^6*7.14^{-11}=1.43*10^{-4}\:\rm m"